OpenStudy (anonymous):

please help solve this pre cal problem! ln8-ln(x-9)=4

OpenStudy (anonymous):

& please show steps so I can learn :)

OpenStudy (jdoe0001):

\(\bf ln(8)-ln(x-9)=4\implies ln\left(\cfrac{8}{x-9}\right)=4\implies log_e\left(\cfrac{8}{x-9}\right)=4 \\ \quad \\ \textit{log cancellation rule of }\large a^{log_ax}=x\qquad thus \\ \quad \\ \large log_e\left(\cfrac{8}{x-9}\right)=4\implies {\color{red}{ e}}^{log_e\left(\frac{8}{x-9}\right)}={\color{red}{ e}}^4\implies \cfrac{8}{x-9}=e^4\)

OpenStudy (jdoe0001):

surely you can take it from there

OpenStudy (anonymous):

from there would i multiply both sides by x-9?

OpenStudy (jdoe0001):

yeap, cross-multiply and solve for "x"

OpenStudy (anonymous):

thank you so much!! :)

OpenStudy (jdoe0001):

yw